# Predicting potentially unfair clauses in Chilean terms of services with natural language processing

### v1

@@ 5.2 Classification
In our proposed approach for the corpus, classification follows on detection. This chapter demonstrates the capabilities of a diverse set of models to classify potentially abusive clauses for each of the three subsets Illegal Classification, Dark Classification and Gray Classification, utilizing fine-tuning or few-shot learning.
@@
Table 6: Experimental results for the classification tasks. We group fine-tuned LLMs into models, that were pre-trained on English, or those, that were pre-trained on Spanish or more languages. We group LLMs for few-shot learning into commercial or open models. Few-shot evaluations are performed with $n=1$. In addition, we include results for $n=5$ and $n=10$ for GPT-4o and -mini. The other LLMs did not benefit from higher $n$. The SVM with TF-IDF features serves as baseline. For convenience, we print the overall best results in bold, and underline the results that are best per group.
The Illegal Classification dataset makes up about 45% of the corpus and consists of classes with many examples each. In our experiments, we observe that the largest fine-tuned model XLM-RoBERTa-large shows the highest $\mu$-F1 score of $78\%$ and m-F1 score of $70\%$, and significantly outperforms few-shot learners. This may be due to the circumstance, that Illegal Clauses form a large enough training dataset to generalize well to unseen examples, and that Illegal Clauses infringe specific laws using specific wording, such as references to non-Chilean jurisdictions, that aids the classifier model. From the remaining fine-tuned models, within the group of English-language models, DeBERTaV3 leads significantly in m-F1 score by 4%, demonstrating benefits of its architectural advances compared to the baseline transformers like BERT. However, the broad Spanish-language pre-training of BETO uncased, and to a lesser extend of RoBERTalex, compensates for DeBERTaV3’s architectural advantages. Overall, the fine-tuned Spanish-language or multilingual models perform on par or better than English-language models.
@@
The Dark Classification clauses are manifestly unfair clauses defined by Article 16 LPC. Here, difficulties may arise from clause reformulation and other intra-class variability. In addition, the class imbalance combined with a multi-label classification may proof difficult. We observe different types of approaches leading in $m$-F1 or $\mu$-F1 scores. In $\mu$-F1, the highly specialized Spanish-language models BETO and RoBERTalex lead by up to 4% over the the best English-language model DeBERTaV3 and 6% over the best few-shot learners GPT-4o/-mini. However, while they dominate on the $\mu$-F1 score, the GPT-variants and several open LLMs perform stronger in the $m$-F1 metric. For GPT-4o-mini, this lead is $5$ to $10\%$ $m$-F1 Score compared to the best fine-tuned models. The open LLMs Gemma2 27B and both Qwen2.5 variants are competitive in $m$-F1 scores but disappoint in the $\mu$-F1 metric. These interesting differences between learning approaches may be due to fewer available training data for fine-tuning, while the GPT models can utilize their world knowledge to predict rare classes from only a few available examples.
@@
In summary, our study shows that ML can significantly aid in the classification of potentially unfair clauses on the Chilean ToS corpus. The differences in performance between the Illegal, Dark and Gray Classification indicate the importance of a suitable model selection and knowledge of contextual facts, given the datasets’ high levels of ambiguity (see Fig. 1) and diverse training data sizes (see Tab. 3). Leveraging language- and domain-specific pre-training, and modern architectures and training methods, as well as in-context learning with large shot counts, can significantly improve classification.
@@ 5.3 Error Analysis
This section conducts a more profound performance analysis to show the difficulties in the proposed datasets. To this end, we present class-wise precision and recall results. Figure 7 shows classification reports for the three datasets. We analyze the fine-tuned models with highest m-F1 score per dataset: BETO uncased for Illegal Classification, XLM-RoBERTa-large for Dark Classification, and RoBERTalex for Gray Classification.
On Illegal Classification, the report indicates that especially "lpc jus" is hard to predict. This may be due to the diversity of this case of potentially unfair clauses. It includes a variety of ways that contracts limit access to justice, such as the ability to access justice directly, or infringements upon time limits, the burden of proof, presumptions, or the court competencies. We record the second-worst scores for the class "lpc", that acts as a residual class and contains less specific clauses. In contrast, the remaining labels "na", "lpc pro", "rc" and "lpc int" are more specific and thus more reliably classified.
On Dark Classification, XLM-RoBERTa-large is severely challenged. The dataset is highly imbalanced. The two classes "ch" and "ter" are extremely rare, and thus, hard to classify correctly. Furthermore, the clauses are often annotated with multiple labels, e.g., as both one-sided (er) and limiting the company’s liability (ltd). In consequence, the model overfits to the much more frequent "ltd" class (precision 85%, recall 86%) and does not recall "er" well (recall 33%).
On Gray Classification, samples are not only highly ambiguous but also annotated with up to three labels per sample. The initially surprising result of perfect precision scores for four of the eight classes ("des det", "des def", "des inf" and "des lic") can be explained by their low rate of overlap with other classes. Only "des det", that also scores the lowest recall score of the eight classes, co-occurs often with other classes. The three most frequently classes "des risk", "bfe" and "des reser" co-occur at a high rate, which leads to higher confusion that expresses itself in the lower scores.
@@ 6 Discussion
ML is becoming an increasingly potent tool to process the large volume of information generated within the legal domain [31]. Especially the family of transformer-based language models has come into widespread use with models ranging from millions [18, 14, 12, 10, 26] to billions [47, 63, 22, 30, 41, 42] of parameters. However, the main line of research has long been focused on English-language models due to the high availability of datasets [18, 39, 33, 7]. With the success of these methods, multilingual [14, 18] and language specific models, e.g., the Spanish-language model BETO [10], and domain-specific models for legal problems, e.g, the English-only Legal-BERT [12] or recently Lawma [24], gained popularity. Even though such language- and domain-specific models may perform better in their designated tasks, they are harder to develop and to evaluate. For example, the Spanish legalese model RoBERTalex [26] has not yet been evaluated on suitable tasks due to a lack of high-quality domain-specific datasets.
[31] recognize that free-access legal resources are often limited, and the lack of datasets is a challenge in legal NLP. We find this issue exacerbated by the focus on the English language and major jurisdictions, such as the European Union or the US. In the realm of consumer contract analysis, we mainly find few and mainly European law datasets [6, 40, 52, 21, 23]. Instead of new datasets, new publications often appear as curated dataset collections [13, 29, 25, 46]. While these help drive legal NLP’s method development  [29, 3] their scope often remains limited. In this publication, we address the critical gap of non-English-language legal datasets that are not from the usual major jurisdictions, and present a novel Spanish-language legal dataset of Chilean law. In addition, we propose a large set of strong baseline methods to measure future developments against. This work sets the stage for future research in the legal context.
In the rest of this section, we discuss transformer-based models in Sec. 6.1, relevant datasets in Sec. 6.2 and conclude the discussion with an outlook in Sec. 6.3.
@@ 6.1 Transformer-based Models
With the introduction of the Attention mechanism [61], a new wave of model architectures emerged, with BERT [18] or the improved RoBERTa [39] quickly adopted as baseline methods for language encoding. M-BERT [18] introduced BERT’s multilingual capabilities, and soon after RoBERTa and XLM-RoBERTa [14] followed suit. Rapid architectural advances ensued, among others, Longformer [7] increased the limiting short sequence length of 512 tokens to 4096 and DeBERTaV3 [33] introduced disentangled attention that improved contextual understanding. [31] provide a more complete systematic overview of transformer models with a focus on legal tasks. Most importantly, they also include an overview of non-English BERT models, such as BETO [10] and RoBERTalex [26], and highlight the challenge of training domain- and language-specific BERT variants like the English-only Legal-BERT [12].
@@
Domain-specific pre-training is another promising approach to transfer-learning. [12] introduce Legal-BERT, a model pre-trained on English-language legal corpora. When compared with BERT on legal tasks, it manages to outperform it. For Spanish legal texts, [26] proposed RoBERTalex. This model was trained on a large collection of Spanish legal corpora. However, its evaluation by [27] is only on general Named Entity Recognition or classification tasks where M-BERT outperforms it significantly. While the optimal pre-training strategy may vary case-to-case, it is recommended to pre-train on domain-specific corpora [12, 23]. However, the historic lack of domain-specific evaluation datasets [31] made this difficult to confirm for Spanish-language legal tasks. We investigate this by evaluating multilingual, language-specific, domain-specific and language- and domain-specific models. We contribute the first evaluation of RoBERTalex on a Spanish-language legal classification dataset. In contrast to the general recommendation to pre-train on a domain-specific corpora [12, 23] we actually see multilingual models like XLM-RoBERTa going head-to-head with RoBERTalex when considering the same number of trainable parameters, depending on the ambiguity of the dataset.
We even see that the more ambiguous tasks Dark Classification and Gray Classification are better addressed by few-shot learning via multilingual LLMs altogether. The training corpora of LLMs usually consists of more languages than just English, with French, German, Chinese and Spanish being the most common minority languages [62]. In variations, this applies to the models that we evaluated in our study. This helps explain the multilingual capabilities of LLMs. Coupled with cross-language transfer [45], this leads to strong generalization capabilities in the low-resource few-shot learning setting, especially with the Gray Classification dataset.
@@ 6.2 Legal Datasets
Free-access legal resources are a limiting factor in the legal NLP domain [31]. This is especially true for non-English languages and exacerbated by the focus on major jurisdictions. This section provides a concise overview of available datasets for consumer contract analysis in the English-language, the multilingual and especially the Spanish-language legal domain.
@@
[40] propose the Unfair Terms of Service dataset with 50 European consumer contracts in English annotated with 8 types of potentially abusive clauses. The authors propose models such as SVMs, Hidden-Markov Models (HMMs), Convolutional Neural Networks (CNNs) or Long-short Term Memory (LSTM) networks to detect and to classify the clauses. [52] extend the initial dataset to a total of 100 consumer contracts, but reduce the annotations to only 5 types of potentially abusive clauses. Their propose method provides explanations by augmenting the classifier with memory, that enriches the data available to the predictor and that provides insights into the networks’ reasoning. This line of investigation spawned numerous methodological developments, such as [3] and LexGLUE [13], that perform fine-tuning studies with more recent model architectures.
@@
The early example of a multilingual contract analysis dataset by [21] contains 25 contracts in four languages (English, German, Italian and Polish) and annotates 9 types of potentially unfair clauses from European consumer law, focusing on sentence-level annotation transfer across the four languages using ELMo [49] sentence embeddings coupled with Dynamic Time Warping [54]. Recently,  [28] evaluate a set of methods to transfer annotations between contracts provided in different languages. They evaluate ELMo and BERT embeddings, besides the traditional Bag-of-words approach. The authors conduct their experiments on an extended multilingual Unfair ToS [21] with 50 contracts per language (200 in total).
@@
Several publications analyze Terms of Services in European member states. [9] propose a new annotation scheme for classifying relevant clauses in Italian Terms and Conditions contracts. However, their work does not include the legal analysis whether the clause are potentially abusive. [6] propose the AGB-DE dataset of 3,764 annotated clauses of German consumer contracts and compare the classification performance of a baseline SVM, fine-tuned language models and GPT-3.5. Fine-tuned models achieve higher precision, GPT-3.5 a higher recall, which is in line with our results. The authors highlight the difficulty of classification due to the ambiguity of complex clauses. [23] publish a Polish-language datasets of 24.000 clauses labeled as safe or abusive for binary classification. Interestingly, their ML method casts the abusiveness of clauses as a problem of similarity-based information retrieval. If a clause is similar to other clauses, that were previously annotated as abusive by legal experts, it may also be abusive. Their method determines similarity between clauses based on distances between neural embeddings, conceptually similar to work by [52].
@@
As an example of Chinese legal tasks, we include the work of [25] in our discussion. The authors propose diverse legal tasks in the Chinese language, categorized in the dimensions of memorization of knowledge, understanding of texts and applied reasoning. They evaluate, among others, multi-label classification, and conclude that GPT-4 remains the best-performing LLM even compared with open multi-lingual and Chinese-oriented LLMs, and even LLMs fine-tuned on Chinese legal corpora.
@@
Recently, Spanish-language legal datasets were published.  [2] published a large Spanish-language legal datasets of consumer Terms of Services contracts. However, their focus is on information retrieval of topics instead of potentially abusive clauses. [20] generate four Spanish legal datasets from the MAPA project for the entity detection task. Similarly, [46] propose the LEXTREME dataset collection that contains two datasets with Spanish law tasks from the context of the European Union, namely MultiEURLEX for document classification and MAPA for Named Entity Recognition. To the best of our knowledge, we contribute the first dataset for Spanish-language (multi-label) clause classification, and the first legal dataset for Chilean law.

### v2

@@ 5.2 Classification
In our proposed approach for the corpus, classification follows on detection. This chapter demonstrates the capabilities of a diverse set of models to classify potentially abusive clauses for each of the three subsets Illegal Classification, Dark Classification, and Gray Classification, utilizing fine-tuning or few-shot learning.
@@
Table 6: Experimental results for the classification tasks. We group fine-tuned LLMs into models, that were pre-trained in English, or those, that were pre-trained in Spanish or more languages. We group LLMs for few-shot learning into commercial or open models. Few-shot evaluations are performed with $n=1$. In addition, we include results for $n=5$ and $n=10$ for GPT-4o and -mini. The other LLMs did not benefit from higher $n$. The SVM with TF-IDF features serves as the baseline. For convenience, we print the overall best results in bold and underline the results that are best per group.
The Illegal Classification dataset makes up about 45% of the corpus and consists of classes with many examples each. In our experiments, we observe that the largest fine-tuned model XLM-RoBERTa-large shows the highest $\mu$-F1 score of $78\%$ and m-F1 score of $70\%$, and significantly outperforms few-shot learners. This may be due to the circumstance, that Illegal Clauses form a large enough training dataset to generalize well to unseen examples, and that Illegal Clauses infringe specific laws using specific wording, such as references to non-Chilean jurisdictions, that aids the classifier model. From the remaining fine-tuned models, within the group of English-language models, DeBERTaV3 leads significantly in the m-F1 score by 4%, demonstrating the benefits of its architectural advances compared to the baseline transformers like BERT. However, the broad Spanish-language pre-training of BETO uncased, and to a lesser extent of RoBERTalex, compensates for DeBERTaV3’s architectural advantages. Overall, the fine-tuned Spanish-language or multilingual models perform on par or better than English-language models.
@@
The Dark Classification clauses are manifestly abusive clauses defined by Article 16 LPC. Here, difficulties may arise from clause reformulation and other intra-class variability. In addition, the class imbalance combined with a multi-label classification may prove difficult. We observe different types of approaches leading in $m$-F1 or $\mu$-F1 scores. In $\mu$-F1, the highly specialized Spanish-language models BETO and RoBERTalex lead by up to 4% over the best English-language model DeBERTaV3 and 6% over the best few-shot learners GPT-4o/-mini. However, while they dominate on the $\mu$-F1 score, the GPT-variants and several open LLMs perform stronger in the $m$-F1 metric. For GPT-4o-mini, this lead is $5$ to $10\%$ $m$-F1 Score compared to the best fine-tuned models. The open LLMs Gemma2 27B and both Qwen2.5 variants are competitive in $m$-F1 scores but disappoint in the $\mu$-F1 metric. These interesting differences between learning approaches may be due to fewer available training data for fine-tuning, while the GPT models can utilize their world knowledge to predict rare classes from only a few available examples.
@@
In summary, our study shows that ML can significantly aid in the classification of potentially abusive clauses on the Chilean ToS corpus. The differences in performance between the Illegal, Dark, and Gray Classification indicate the importance of a suitable model selection and knowledge of contextual facts, given the datasets’ high levels of ambiguity (see Fig. 1) and diverse training data sizes (see Tab. 3). Leveraging language- and domain-specific pre-training, and modern architectures and training methods, as well as in-context learning with large shot counts, can significantly improve classification.
@@ 5.3 Error Analysis
This section conducts a more profound performance analysis to show the difficulties in the proposed datasets. To this end, we present class-wise precision and recall results. Figure 7 shows classification reports for the three datasets. We analyze the fine-tuned models with the highest m-F1 score per dataset: BETO uncased for Illegal Classification, XLM-RoBERTa-large for Dark Classification, and RoBERTalex for Gray Classification.
@@
On Illegal Classification, the report indicates that especially "lpc jus" is hard to predict for BETO. This may be due to the diversity of this case of potentially abusive clauses. It includes a variety of ways that contracts limit access to justice, such as the ability to access justice directly, or infringements upon time limits, the burden of proof, presumptions, or the court competencies, e.g., the following clause restricts access to justice in Chile. However, the model predicts this instance incorrectly as the more specific "lpc pro":
You and Nintendo agree that the state or federal courts in King County, Washington have exclusive jurisdiction over any appeal of an arbitration award and over any claim between the parties that has not been submitted to arbitration. Disputes between the parties shall be governed by this Agreement and the applicable laws of the state of Washington and the United States, without giving effect to principles of conflicts of laws that would permit the application of the law of another jurisdiction. [^12] Translated from Spanish using DeepL.com: ”Usted y Nintendo acuerdan que los tribunales estatales o federales del condado de King, Washington, tienen jurisdicción exclusiva sobre cualquier apelación de un laudo arbitral y sobre cualquier demanda entre las partes que no se haya sometido a arbitraje. Los conflictos entre las partes se regirán por este Acuerdo y las leyes aplicables del estado de Washington y de Estados Unidos, sin que se ejecuten principios de conflicto de leyes que permitan la aplicación de la ley de otra jurisdicción.”
This example emphasizes the difficulty of predicting the distinction between a specific alternation of the LPC’s jurisdiction rules ("lpc pro") and the more general limitation of access to justice ("lpc jus").
We record the second-worst scores for the class "lpc", which acts as a residual class and contains fewer specific clauses. In contrast, the remaining labels "na", "lpc pro", "rc" and "lpc int" are more specific and thus more reliably classified.
On Dark Classification, XLM-RoBERTa-large is severely challenged. The dataset is highly imbalanced. The two classes "ch" and "ter" are extremely rare, and thus, hard to learn correctly. The following example shows a clause that allows the unilateral modification of the price of a service:
The fees we charge sellers for using our Services to sell products and services are set forth on our Selling Rates pages. We may change our selling fees from time to time. Such changes will be posted on the eBay site fourteen (14) days in advance. Prior notice is not required for temporary promotions or any changes that result in a reduction of fees. [^13] Translated from Spanish using DeepL.com: ”Las tarifas que les cobramos a los vendedores por usar nuestros Servicios para vender productos y servicios se indican en nuestras páginas de Tarifas de venta. Podemos cambiar nuestras tarifas de venta periódicamente. Dichos cambios se publicarán en el sitio de eBay con catorce (14) días de antelación. No se requiere una notificación previa para promociones temporales o cualquier cambio que tenga como consecuencia una reducción de las tarifas.”
The model is not leveraging the relevant keywords for the following exemplary clause of the rare class "ch", such as "change", "fees" or "Selling Rates" in their respective Spanish original. Instead, the model predicts "cr", which refers to a unilateral change of the ToS without users being able to challenge or reject the changes.
Furthermore, the clauses are often annotated with multiple labels, e.g., as both one-sided (er) and limiting the company’s liability (ltd). In consequence, the model overfits to the much more frequent "ltd" class (precision 85%, recall 86%) and does not recall "er" well (recall 33%).
On Gray Classification, samples are not only highly ambiguous but also annotated with up to three labels per sample. The initially surprising result of perfect precision scores for four of the eight classes ("des det", "des def", "des inf" and "des lic") can be explained by their low rate of overlap with other classes. Only "des det", which also scores the lowest recall score of the eight classes, often co-occurs with other classes. The three most frequent classes "des risk", "bfe" and "des reser" co-occur at a high rate, which leads to higher confusion that expresses itself in the lower scores. An example of this issue is the following "des det" clause, which affects the processes for resolving conflicts:
In the event that the User breaches a law or the Terms and Conditions, we may warn, suspend, restrict or temporarily or permanently disable his/her account, without prejudice to other sanctions established in the particular rules of use of Mercado Libre’s services. [^14] Translated from Spanish using DeepL.com: ”En caso que la Persona Usuaria incumpliera una ley o los Términos y Condiciones, podremos advertir, suspender, restringir o inhabilitar temporal o definitivamente su cuenta, sin perjuicio de otras sanciones que se establezcan en las reglas de uso particulares de los servicios de Mercado Libre.”
RoBETalex incorrectly classifies this as "des reser", which allows the provider to modify the contract at their discretion. This error may be explained due to the relatively high co-occurrence of both classes in the dataset, see Fig. 1 for reference.
@@ 6 Discussion
ML is becoming an increasingly potent tool to process the large volume of information generated within the legal domain [31]. Especially the family of transformer-based language models has come into widespread use with models ranging from millions [18, 14, 12, 10, 26] to billions [48, 64, 22, 30, 41, 42] of parameters. However, the main line of research has long been focused on English-language models due to the high availability of datasets [18, 39, 33, 7]. With the success of these methods, multilingual [14, 18] and language-specific models, e.g., the Spanish-language model BETO [10], and domain-specific models for legal problems, e.g, the English-only Legal-BERT [12] or recently Lawma [24], gained popularity. Even though such language- and domain-specific models may perform better in their designated tasks, they are harder to develop and evaluate. For example, the Spanish legalese model RoBERTalex [26] has not yet been evaluated on suitable tasks due to a lack of high-quality domain-specific datasets.
[31] recognize that free-access legal resources are often limited, and the lack of datasets is a challenge in legal NLP. We find this issue exacerbated by the focus on the English language and major jurisdictions, such as the European Union or the US. In the realm of consumer contract analysis, we mainly find few and mainly European law datasets [6, 40, 53, 21, 23]. Instead of new datasets, new publications often appear as curated dataset collections [13, 29, 25, 47]. While these help drive legal NLP’s method development [29, 3] their scope often remains limited. In this publication, we address the critical gap of non-English-language legal datasets that are not from the usual major jurisdictions and present a novel Spanish-language legal dataset of Chilean law. In addition, we propose a large set of strong baseline methods against which to measure future developments. This work sets the stage for future research in the legal context.
In the rest of this section, we discuss transformer-based models in Sec. 6.1, relevant datasets in Sec. 6.2, and conclude the discussion with an outlook in Sec. 6.3.
@@ 6.1 Transformer-based Models
With the introduction of the Attention mechanism [62], a new wave of model architectures emerged, with BERT [18] or the improved RoBERTa [39] quickly adopted as baseline methods for language encoding. M-BERT [18] introduced BERT’s multilingual capabilities, and soon after RoBERTa and XLM-RoBERTa [14] followed suit. Rapid architectural advances ensued, among others, Longformer [7] increased the limiting short sequence length of 512 tokens to 4096 and DeBERTaV3 [33] introduced disentangled attention that improved contextual understanding. [31] provide a complete systematic overview of transformer models focusing on legal tasks. Most importantly, they also include an overview of non-English BERT models, such as BETO [10] and RoBERTalex [26], and highlight the challenge of training domain- and language-specific BERT variants like the English-only Legal-BERT [12].
@@
Domain-specific pre-training is another promising approach to transfer learning. [12] introduce Legal-BERT, a model pre-trained on English-language legal corpora. When compared with BERT on legal tasks, it manages to outperform it. For Spanish legal texts, [26] proposed RoBERTalex. This model was trained on a large collection of Spanish legal corpora. However, its evaluation by [27] is only on general Named Entity Recognition or classification tasks where M-BERT outperforms it significantly. While the optimal pre-training strategy may vary case-to-case, it is recommended to pre-train on domain-specific corpora [12, 23]. However, the historic lack of domain-specific evaluation datasets [31] made this difficult to confirm for Spanish-language legal tasks. We investigate this by evaluating multilingual, language-specific, domain-specific, and language- and domain-specific models. We contribute the first evaluation of RoBERTalex on a Spanish-language legal classification dataset. In contrast to the general recommendation to pre-train on a domain-specific corpora [12, 23] we actually see multilingual models like XLM-RoBERTa going head-to-head with RoBERTalex when considering the same number of trainable parameters, depending on the ambiguity of the dataset.
We even see that the more ambiguous tasks Dark Classification and Gray Classification are better addressed by few-shot learning via multilingual LLMs altogether. The training corpora of LLMs usually consist of more languages than just English, with French, German, Chinese, and Spanish being the most common minority languages [63]. In variations, this applies to the models that we evaluated in our study. This helps explain the multilingual capabilities of LLMs. Coupled with cross-language transfer [46], this leads to strong generalization capabilities in the low-resource few-shot learning setting, especially with the Gray Classification dataset.
@@ 6.2 Legal Datasets
Free-access legal resources are a limiting factor in the legal NLP domain [31]. This is especially true for non-English languages and exacerbated by the focus on major jurisdictions. This section provides a concise overview of available datasets for consumer contract analysis in the English language, the multilingual, and especially the Spanish-language legal domain.
@@
[40] propose the Unfair Terms of Service dataset with 50 European consumer contracts in English annotated with 8 types of potentially abusive clauses. The authors propose models such as SVMs, Hidden-Markov Models (HMMs), Convolutional Neural Networks (CNNs), or Long-short Term Memory (LSTM) networks to detect and classify the clauses. [53] extend the initial dataset to a total of 100 consumer contracts, but reduce the annotations to only 5 types of potentially abusive clauses. Their proposed method provides explanations by augmenting the classifier with memory, which enriches the data available to the predictor and provides insights into the networks’ reasoning. This line of investigation spawned numerous methodological developments, such as [3] and LexGLUE [13], that perform fine-tuning studies with more recent model architectures.
@@
The early example of a multilingual contract analysis dataset by [21] contains 25 contracts in four languages (English, German, Italian and Polish) and annotates 9 types of potentially unfair clauses from European consumer law, focusing on sentence-level annotation transfer across the four languages using ELMo [50] sentence embeddings coupled with Dynamic Time Warping [55]. Recently,  [28] evaluated a set of methods to transfer annotations between contracts provided in different languages. They evaluate ELMo and BERT embeddings, besides the traditional Bag-of-words approach. The authors conduct their experiments on an extended multilingual Unfair ToS [21] with 50 contracts per language (200 in total).
@@
Several publications analyze the Terms of Services in European member states. [5] propose a new annotation scheme for classifying relevant clauses in Italian Terms and Conditions contracts. However, their work does not include a legal analysis of whether the clauses are potentially abusive. [6] propose the AGB-DE dataset of 3,764 annotated clauses of German consumer contracts and compare the classification performance of a baseline SVM, fine-tuned language models and GPT-3.5. Fine-tuned models achieve higher precision, GPT-3.5 a higher recall, which is in line with our results. The authors highlight the difficulty of classification due to the ambiguity of complex clauses. [23] publish a Polish-language dataset of 24.000 clauses labeled as safe or abusive for binary classification. Interestingly, their ML method casts the abusiveness of clauses as a problem of similarity-based information retrieval. If a clause is similar to other clauses, that were previously annotated as abusive by legal experts, it may also be abusive. Their method determines the similarity between clauses based on distances between neural embeddings, conceptually similar to work by [53].
@@
As an example of Chinese legal tasks, we include the work of [25] in our discussion. The authors propose diverse legal tasks in the Chinese language, categorized in the dimensions of memorization of knowledge, understanding of texts, and applied reasoning. They evaluate, among others, multi-label classification, and conclude that GPT-4 remains the best-performing LLM even compared with open multi-lingual and Chinese-oriented LLMs, and even LLMs fine-tuned on Chinese legal corpora.
@@
Recently, a few Spanish-language legal datasets were published.  [2] published a large Spanish-language legal dataset of consumer Terms of Services contracts. However, their focus is on information retrieval of topics instead of potentially abusive clauses. [20] generate four Spanish legal datasets from the MAPA project for the entity detection task. Similarly, [47] propose the LEXTREME dataset collection that contains two datasets with Spanish law tasks from the context of the European Union, namely MultiEURLEX for document classification and MAPA for Named Entity Recognition. To the best of our knowledge, we contribute the first dataset for Spanish-language (multi-label) clause classification, and the first legal dataset for Chilean law.